OpenStudy (anonymous):

Help solve: (8/(c^2-4))+(2/(c^2-5c+6))

OpenStudy (anonymous):

I know that it is (8/(c-2)(c+2))+(2/(c-2)(c-3))

OpenStudy (anonymous):

By 'solve', would that be : \[\frac{8}{c^{2}-4}+ \frac{2}{c^2-5c+6}=0\]?? If so then factorize a bit:\[\frac{8}{(c+2)(c-2)}+\frac{2}{(c-3)(c-2)}=0\]so this initially tells you one cannot have c =-2, 2 or +3 as one or more denominators is zero. Now multiply both sides by (c+2)(c-2)(c-3) yielding:\[8(c-3)+2(c+2)=0\]\[8c-24+2c+4=0\]\[10c=+20\]or c=2 but that was disallowed by the presumed problem specification.